OpenStudy (anonymous):

-5x-9y=-8 Find 3 Ordered Pairs

OpenStudy (anonymous):

can you solve for y

OpenStudy (mayankdevnani):

value of x=0

OpenStudy (mayankdevnani):

-9y=-8 y=8/9

OpenStudy (anonymous):

-5(0)-9y=-8

OpenStudy (mayankdevnani):

first order pair (0,8/9)

OpenStudy (anonymous):

Right But How Will I Graph (0,8/9)

OpenStudy (mayankdevnani):

8/9=0.9

OpenStudy (mayankdevnani):

something

OpenStudy (anonymous):

you chose 0 for x so you get (0,y) plug 0 in the equation and solve for y and you get (0,8/9)

OpenStudy (anonymous):

Okay But How Will I Graph (0,8/9)

OpenStudy (mayankdevnani):

 ok

OpenStudy (mayankdevnani):

got it

OpenStudy (anonymous):

What Number Is It On

OpenStudy (mayankdevnani):

0.9